# Online Multi-Task Offloading for Semantic-Aware Edge Computing Systems

### v2

@@ III-B Semantic-aware MAPPO Algorithm
4:     Calculate $T$ step GAE $A^{GAE}_{1},...,A^{GAE}_{T}$ using $\pi_{\boldsymbol{\theta^{\prime}}}$ and $\boldsymbol{\phi}$;
5:     for epoches do
@@ IV Simulation Results
IV-A Implementation Details
Three tasks are considered: text classification, image classification, and VQA, corresponding to text, image, and multi-modal tasks, respectively. For text classification, we use the SST-2 dataset, for image classification, the CIFAR-10 dataset, and for VQA, the VQAv2 dataset is adopted. The text semantic extraction network utilizes an advanced BERT-based [45] text embedding model, while the image semantic extraction network is based on the Vision Transformer (ViT) [46]. For VQA, the encoder architecture employs BERT for linguistic processing and ViT for visual feature extraction, with cross-modal fusion and answer generation performed at the receiver. Training is conducted using the AdamW optimizer, with a learning rate of $3\times 10^{-5}$, a batch size of 50, and a weight decay of $1\times 10^{-4}$. All experiments were conducted on a platform equipped with an NVIDIA RTX 4090 and a CPU Intel i9-13900K @5.8 GHz. When the training of all semantic transmission models is complete, we can calculate the GPU consumption for basic semantic extraction $l_{n}^{SE}$ and the GPU consumption $l^{S}$ for execution at the ES.
We consider scenarios with four UEs and orthogonal sub-carriers, where each UE has a queue length of 20, containing tasks from different task types. Additional details regarding the wireless environment, GPU computational frequency, and the proposed semantic-aware MAPPO algorithm training configurations are summarized in Table II. To verify the effectiveness of the semantic-aware multi-task offloading framework, we consider the following benchmarks:
TABLE II: Simulation Parameters
Parameter Value
Number of UEs $N$ 4
@@ IV-A Implementation Details
  • Semantic-unaware MAPPO: The benchmark transmits all data without semantic extraction during offloading, i.e. $\mu_{n}=1$. All other settings are identical to those used in semantic-aware MAPPO.
  • Semantic-aware Dueling Double DQN (D3QN): D3QN can only handle discrete action spaces, so we discretized the action space and used the same environment as semantic-aware MAPPO.
  • Semantic-unaware D3QN: This benchmark transmits all data without semantic extraction during offloading, i.e. $\mu_{n}=1$. All other settings are identical to those used in semantic-aware D3QN.
@@
IV-B QoE Performance and Convergence
1) Convergence Performance: The preference parameters, $\omega_{t}$, $\omega_{e}$, and $\omega_{a}$ are all set to 1/3, indicating that the model has no specific preference among latency, energy consumption, and accuracy. Fig. 6 shows that the proposed MAPPO algorithm converges after approximately 100 episodes and maintains optimal performance. For the D3QN algorithm, the replay buffer size is set to 30000. Fig. 7 shows that the D3QN algorithm finally converges after 2000 episodes. The proposed MAPPO algorithm directly optimizes cumulative returns through policy gradient methods, while employing a trust region mechanism to constrain policy update magnitudes, thereby preventing training instability caused by abrupt policy changes. In contrast, the D3QN algorithm relies on TD learning and demonstrates lower learning efficiency in environments with sparse rewards or delayed feedback. Furthermore, to enhance the sampling efficiency of on-policy algorithms, the proposed MAPPO algorithm employs multi-threaded asynchronous sampling and completes state collection and synchronization during network updates. Consequently, the proposed MAPPO algorithm achieves faster convergence with greater stability compared to D3QN.
@@ IV-B QoE Performance and Convergence
2) QoE Performance: Fig. (8) reveals the relationship between QoE and the sub-carrier bandwidth for UEs. All schemes are initially trained at a sub-carrier bandwidth of 10 MHz and then tested under various bandwidth conditions. It is observed that the proposed semantic-aware MAPPO algorithm achieves the best performance across all bandwidth conditions. Comparative evaluations demonstrate 12.68% QoE improvements versus semantic-unaware MAPPO and 14.48% superiority versus semantic-unaware D3QN. Semantic-unaware methods fix $\mu_{n}=1$ to transmit all data without compression, increasing transmission and execution latency, consequently degrading QoE. The D3QN algorithm requires discrete action space approximation, which may introduce quantization errors or reduce sampling efficiency. In contrast, the proposed MAPPO algorithm directly outputs continuous action probability distributions, eliminating the need for discretization and thereby avoiding the curse of dimensionality in high-dimensional action spaces. Consequently, our method achieves superior QoE. Local execution is solely related to the local computation frequency. When calculating QoE, the local computation frequency is fixed, meaning we use the local execution’s latency, energy consumption, and accuracy to normalize the latency, energy, and accuracy obtained from other methods. Thus, the QoE of local execution is fixed at 0.5.
Figure 8: QoE with different sub-carrier bandwidths.
@@
Figure 11: QoE with different semantic extraction factor $\mu_{n}$.
Fig. (9) describes how QoE varies with channel noise intensity. All schemes are first trained under the noise intensity of 10 mW and then tested at different noise intensities. Similarly, the proposed semantic-aware MAPPO algorithm performs better than all other benchmarks. The semantic-aware algorithm dynamically selects optimal semantic extraction factors based on real-time noise intensity and channel gain, maintaining stable QoE without significant degradation even under increasing noise levels. In contrast, semantic-unaware schemes lack adaptive semantic symbol quantity regulation, resulting in substantial performance deterioration. Fig. (10) illustrates the QoE trend versus the number of users. As the number of users increases, the server must handle more computation offloading requests, leading to elevated processing latency and energy consumption, which collectively degrade overall QoE. Semantic-unaware approaches cannot leverage semantic extraction to reduce uplink latency during user scaling, ultimately forcing all users to execute tasks locally. Consequently, its QoE converges to the local computing baseline. Fig. (11) illustrates the QoE variation versus $\mu_{n}$. While trained at $\mu_{n}=0.1$, the system is evaluated across varying $\mu_{n}$. Results demonstrate that when users have fewer selectable semantic extraction factors, the inability to utilize smaller compression factors for data size reduction significantly impairs channel adaptation under poor conditions, ultimately degrading QoE.
TABLE III: Performance under different preferences
@@
To further investigate the contributions of latency, energy consumption, and task performance to QoE, Fig. (12) illustrates the execution latency, energy consumption, and task accuracy. All metrics are measured under the following conditions: 10 MHz bandwidth, 2 mW noise power, and 4 users. Local task execution demonstrates longer latency and higher energy consumption, yet achieves optimal task performance. This occurs because local processing avoids performance degradation induced by semantic compression. However, the prolonged execution time incurs significant energy overhead. Semantic-unaware approaches exhibit inferior latency performance because of the transmission of raw task data without semantic extraction, resulting in extended transmission delays. However, server-side processing of offloaded tasks requires minimal time, yielding lower energy consumption compared to local execution. Owing to wireless channel effects during task offloading, semantic-unaware methods show slightly reduced task accuracy relative to local processing. The semantic-aware approach dynamically selects optimal semantic extraction factors based on real-time channel conditions, computational load, and transmission requirements, achieving the best trade-off among latency, energy efficiency, and task accuracy. Consequently, it delivers the highest QoE.
Fig. (13) presents the latency, energy consumption, and task performance across different task queue configurations. The system was initially trained on hybrid task scenarios and subsequently evaluated under four distinct test conditions: hybrid tasks, text classification, image classification, and VQA. The results demonstrate our algorithm’s strong adaptability to diverse task types. Specifically, text classification tasks exhibit the lowest latency and energy consumption while maintaining high accuracy, attributable to their smaller data volume and computational complexity. In contrast, image classification and VQA tasks require substantially more resources, resulting in higher energy expenditure and longer processing times. The hybrid task shows performance metrics that approximately equal the arithmetic mean of individual task types, confirming the algorithm’s robustness. The significant performance variations observed across task categories underscore the necessity of employing locally normalized QoE metrics for fair evaluation.
@@
IV-C QoE Preference
We can adjust the model to achieve desired results by modifying QoE preferences for different metrics. Table III shows the performance in execution latency, energy consumption, and task accuracy under varying preference settings. Specifically, we adjust the relative magnitude of $\omega_{t}$, $\omega_{e}$, and $\omega_{a}$ to set preferences for different metrics and retrain to obtain the corresponding model for each. 1) For the delay-preference model, we set $\omega_{t}=1$, $\omega_{e}=0$, and $\omega_{a}=0$. The delay-preference model tends to choose a smaller $\mu_{n}$ to reduce the amount of data transmitted. This results in a partial loss of semantic information, causing a decline in task accuracy. 2) For the energy-preference model, we set $\omega_{t}=0$, $\omega_{e}=1$ and $\omega_{a}=0$. The energy-preference model prioritizes reduction of local processing energy by favoring larger semantic extraction factors, achieving significant energy savings at the cost of substantially increased transmission latency. 3) For the accuracy-preference model, we set $\omega_{t}=0$, $\omega_{e}=0$, and $\omega_{a}=1$. The accuracy-preference model aims to maximize the accuracy of all tasks, prompting it to execute locally. Nevertheless, Local task execution exhibits significantly higher latency and energy consumption.
Fig. 14 shows the performance of latency and task accuracy under different weights by adjusting the weights assigned to latency and accuracy. As the weight for latency increases, tasks are offloaded to the ES with more powerful processing capabilities, and there tends to be a preference for a smaller semantic extraction factor, $\mu_{n}$, to maximize the compression rate. This results in a decrease in the accuracy of tasks executed by the ES. Fig. 15 presents the performance of energy consumption and task accuracy under different weights. A significant jump occurs at (0.8, 0.2) when the weight for accuracy increases, as tasks are executed locally, causing a substantial increase in energy consumption and a corresponding rise in accuracy. Fig. 16 describes the performance of energy consumption and latency under varying weights. When the weight for latency increases beyond (0.3, 0.7), latency begins to decrease, and energy consumption increases. This is due to an increase in the number of tasks offloaded to the ES, and users tend to choose a smaller semantic extraction factor, $\mu_{n}$, to reduce data transmission volume. However, a smaller $\mu_{n}$ means that users need more computational resources to extract more compact semantic information.
@@ IV-C QoE Preference
V Conclusion
In this paper, we proposed a new multi-task offloading system and designed a unified quality of experience (QoE) metric to measure overall task offloading performance. Text classification, image classification, and visual question answering (VQA) tasks can be executed locally or offloaded to the edge server after semantic extraction. A semantic-aware multi-agent PPO reinforcement learning algorithm is used to solve the problem of maximizing overall user QoE through the joint optimization of the semantic extraction factor, computing resources, and communication resources. Experimental results show that the proposed method outperforms other benchmarks in overall user QoE. Specifically, the proposed algorithm demonstrates a 12.68% improvement in QoE compared to the semantic-unaware approach. Additionally, the proposed approach can be flexibly adapted to user preferences, fully satisfying latency, energy, and task performance preference requirements.

### v3

@@ IV-B Semantic-aware MAPPO Algorithm
4:   Calculate $T$ step GAE $A^{\mathrm{GAE}}_{1},...,A^{\mathrm{GAE}}_{T}$ using $\pi_{\boldsymbol{\theta^{\prime}}}$ and $\boldsymbol{\phi}$;
5:   for epochs do
@@ V Simulation Results
V-A Implementation Details
Three tasks are considered: text classification, image classification, and VQA, corresponding to text, image, and multi-modal tasks, respectively. For text classification, we use the SST-2 dataset, for image classification, the CIFAR-10 dataset, and for VQA, the VQAv2 dataset is adopted. The text semantic extraction network utilizes an advanced BERT-based [101] text embedding model, while the image semantic extraction network is based on the Vision Transformer (ViT) [102]. For VQA, the encoder architecture employs BERT for linguistic processing and ViT for visual feature extraction, with cross-modal fusion and answer generation performed at the receiver. The text classification model requires 3.72 GFLOPs, the image classification model requires 8.43 GFLOPs, and the VQA model requires 8.31 GFLOPs. Training is conducted using the AdamW optimizer, with a learning rate of $3\times 10^{-5}$, a batch size of 50, and a weight decay of $1\times 10^{-4}$. All experiments were conducted on a platform equipped with an NVIDIA RTX 4090 and a CPU Intel i9-13900K @5.8 GHz. When the training of all semantic transmission models is complete, we can calculate the GPU consumption for basic semantic extraction $l_{n}^{SE}$ and the GPU consumption $l^{S}$ for execution at the ES.
The architecture of MAPPO is detailed in Table III. The actor network takes the state information as input and first extracts intermediate representations through a two-layer multilayer perceptron (MLP) with 256 hidden units per layer. The resulting features are then fed into a multi-layer gated recurrent unit (GRU) network to capture temporal dependencies and generate action-related features. Finally, an additional MLP layer maps these features to the action probability distribution. Similarly, the critic network takes the state information as input and processes it through a two-layer MLP with 256 hidden units to obtain intermediate features. These features are subsequently passed to a multi-layer GRU network to produce value-related representations, which are then mapped to the state value via a final MLP layer.
Each UE has a queue length of 20, containing tasks from different task types. Additional details regarding the wireless environment, GPU computational frequency, and the proposed semantic-aware MAPPO algorithm training configurations are summarized in Table IV.
In our experiments, the model sizes for different modality-specific tasks range from several tens to a few hundred megabytes, while the computational complexity of the models is approximately 3 to 9 GFLOPs. Taking a commonly used edge computing platform, the NVIDIA Jetson Nano, as an example, which provides a peak computing capability of about 470 GFLOPS (Floating Point Operations per Second), the proposed models can be efficiently deployed and executed under the given computational constraints. Even for devices with more limited computing resources, the proposed framework is fully compatible with standard model optimization techniques, including model pruning, weight quantization (e.g., 8-bit or mixed-precision inference), and knowledge distillation. Furthermore, selecting an appropriate semantic compression factor $\mu$ can also be tailored to the computational capabilities of the current device.
To verify the effectiveness of the semantic-aware multi-task offloading framework, we consider the following benchmarks:
TABLE III: Network Architectures of MAPPO Actor and Critic
Network Component Layer Description Output / Notes
Actor Fully-Connected (256 units), ReLU, LayerNorm Intermediate feature
Fully-Connected (256 units), ReLU, LayerNorm Refined feature
Multi-layer GRU Action feature
Fully-Connected Action probability distribution
Critic Fully-Connected (256 units), ReLU, LayerNorm Intermediate feature
Fully-Connected (256 units), ReLU, LayerNorm Refined feature
Multi-layer GRU Value feature
Fully-Connected State value estimate
TABLE IV: Simulation Parameters
Parameters Values
Number of UEs NN 4
@@ V-A Implementation Details
The slope of logistic function $\lambda,\beta,\eta$ $1\times 10^{3}$, 10, 2
  • Semantic-aware MAPPO (discrete): This benchmark discretizes continuous actions in the hybrid action space, while keeping other settings the same as MAPPO.
  • Semantic-unaware MAPPO: The benchmark without semantic extraction during offloading, i.e. $\mu_{n}=1$. Specifically, this scheme relies on conventional compression techniques and does not support adaptive semantic extraction. All other settings are identical to those used in semantic-aware MAPPO.
  • Semantic-aware Dueling Double DQN (D3QN): D3QN can only handle discrete action spaces, so we also discretized the action space and used the same environment as semantic-aware MAPPO.
  • Semantic-unaware D3QN: The benchmark without semantic extraction during offloading, i.e. $\mu_{n}=1$. All other settings are identical to those used in semantic-aware D3QN.
@@
V-B QoE Performance and Convergence
1) Convergence Performance: The preference parameters, $\omega_{t}$, $\omega_{e}$, and $\omega_{a}$ are all set to 1/3, indicating that the model has no specific preference among latency, energy consumption, and accuracy. Fig. 6 shows that the proposed MAPPO algorithm converges after approximately 100 episodes and maintains optimal performance. The policy entropy also gradually converges, indicating that the policy becomes stable during training. For the D3QN algorithm, the replay buffer size is set to 30000.
Fig. 7 shows that the D3QN algorithm finally converges after 2000 episodes. The proposed MAPPO algorithm directly optimizes cumulative returns through policy gradient methods, while employing a trust region mechanism to constrain policy update magnitudes, thereby preventing training instability caused by abrupt policy changes. In contrast, the D3QN algorithm relies on TD learning and demonstrates lower learning efficiency in environments with sparse rewards or delayed feedback. Furthermore, to enhance the sampling efficiency of on-policy algorithms, the proposed MAPPO algorithm employs multi-threaded asynchronous sampling and completes state collection and synchronization during network updates. Consequently, the proposed MAPPO algorithm achieves faster convergence with greater stability compared to D3QN.
@@ V-B QoE Performance and Convergence
2) QoE Performance: Fig. 8 reveals the relationship between QoE and the sub-channel bandwidth for UEs. All schemes are initially trained at a sub-channel bandwidth of 10 MHz and then tested under various bandwidth conditions. It is observed that the proposed semantic-aware MAPPO algorithm achieves the best performance across all bandwidth conditions. Comparative evaluations demonstrate 12.68% QoE improvements versus semantic-unaware MAPPO and 14.48% superiority versus semantic-unaware D3QN. Semantic-unaware methods fix $\mu_{n}=1$ without semantic compression, increasing transmission and execution latency, consequently degrading QoE. The D3QN algorithm requires discrete action space approximation, which may introduce quantization errors or reduce sampling efficiency. The semantic-aware MAPPO (discrete) method adopts the same discrete action space configuration as D3QN. In contrast, the proposed MAPPO algorithm directly outputs continuous action probability distributions, and support hybrid action space, eliminating the need for discretization and thereby avoiding the curse of dimensionality in high-dimensional action spaces. Consequently, our method achieves superior QoE. Local execution is solely related to the local computation frequency. When calculating QoE, the local computation frequency is fixed, meaning we use the local execution’s latency, energy consumption, and accuracy to normalize the latency, energy, and accuracy obtained from other methods. Thus, the QoE of local execution is fixed at 0.5.
Figure 8: QoE with different sub-channel bandwidths.
@@
Figure 11: QoE with different minimum semantic extraction factor $\mu_{n}$.
Fig. 9 describes how QoE varies with channel noise intensity. All schemes are first trained under the noise intensity of 2 mW and then tested at different noise intensities. Similarly, the proposed semantic-aware MAPPO algorithm performs better than all other benchmarks. The semantic-aware algorithm dynamically selects optimal semantic extraction factors based on real-time noise intensity and channel gain, maintaining stable QoE without significant degradation even under increasing noise levels. In contrast, semantic-unaware schemes lack adaptive semantic symbol quantity regulation, resulting in substantial performance deterioration. Fig. 10 illustrates the QoE trend versus the number of users. As the number of users increases, the server must handle more computation offloading requests, leading to elevated processing latency and energy consumption, which collectively degrade overall QoE. Semantic-unaware approaches cannot leverage semantic extraction to reduce uplink latency during user scaling, ultimately forcing all users to execute tasks locally. Consequently, its QoE converges to the local computing baseline. Fig. 11 illustrates the QoE variation versus different minimum $\mu_{n}$. We retrain the MAPPO under different minimum values of $\mu_{n}$. Results demonstrate that when users have fewer selectable semantic extraction factors, the inability to utilize smaller compression factors for data size reduction significantly impairs channel adaptation under poor conditions, ultimately degrading QoE.
TABLE V: Performance under Different Preferences
@@
To further investigate the contributions of latency, energy consumption, and task performance to QoE, Fig. 12 illustrates the execution latency, energy consumption, and task accuracy. All metrics are measured under the following conditions: 10 MHz bandwidth, 2 mW noise power, and 4 users. Local task execution demonstrates longer latency and higher energy consumption, yet achieves optimal task performance. This occurs because local processing avoids performance degradation induced by semantic compression. However, the prolonged execution time incurs significant energy overhead. Semantic-unaware approaches exhibit inferior latency performance because of the transmission of raw task data without semantic extraction, resulting in extended transmission delays. However, server-side processing of offloaded tasks requires minimal time, yielding lower energy consumption compared to local execution. Owing to wireless channel effects during task offloading, semantic-unaware methods show slightly reduced task accuracy relative to local processing. The semantic-aware approach dynamically selects optimal semantic extraction factors based on real-time channel conditions, computational load, and transmission requirements, achieving the best trade-off among latency, energy efficiency, and task accuracy. Consequently, it delivers the highest QoE.
Fig. 13 presents the latency, energy consumption, and task performance across different task queue configurations. The system was initially trained on hybrid task scenarios and subsequently evaluated under four distinct test conditions: hybrid tasks, text classification, image classification, and VQA. The results demonstrate our algorithm’s strong adaptability to diverse task types. Specifically, text classification tasks exhibit the lowest latency and energy consumption while maintaining high accuracy, attributable to their smaller data volume and computational complexity. In contrast, image classification and VQA tasks require substantially more resources, resulting in higher energy expenditure and longer processing times. The hybrid task shows performance metrics that approximately equal the arithmetic mean of individual task types, confirming the algorithm’s robustness. The significant performance variations observed across task categories underscore the necessity of employing locally normalized QoE metrics for fair evaluation.
@@
V-C QoE Preference
We can adjust the model to achieve desired results by modifying QoE preferences for different metrics. Table V shows the performance in execution latency, energy consumption, and task accuracy under varying preference settings. Specifically, we adjust the relative magnitude of $\omega_{t}$, $\omega_{e}$, and $\omega_{a}$ to set preferences for different metrics and retrain to obtain the corresponding model for each. 1) For the delay-preference model, we set $\omega_{t}=1$, $\omega_{e}=0$, and $\omega_{a}=0$. The delay-preference model tends to choose a smaller $\mu_{n}$ to reduce the amount of data transmitted. This results in a partial loss of semantic information, causing a decline in task accuracy. 2) For the energy-preference model, we set $\omega_{t}=0$, $\omega_{e}=1$ and $\omega_{a}=0$. The energy-preference model prioritizes reduction of local processing energy by favoring larger semantic extraction factors, achieving significant energy savings at the cost of substantially increased transmission latency. 3) For the accuracy-preference model, we set $\omega_{t}=0$, $\omega_{e}=0$, and $\omega_{a}=1$. The accuracy-preference model aims to maximize the accuracy of all tasks, prompting it to execute locally. Nevertheless, Local task execution exhibits significantly higher latency and energy consumption.
To investigate the adaptability of the proposed algorithm to different weights of latency, energy consumption, and task performance, we vary the weight settings and retrain the resource allocation model, and then evaluate the corresponding changes in latency, energy consumption, and task performance. Fig. 14 shows the performance of latency and task accuracy under different weights by adjusting the weights assigned to latency and accuracy. As the weight for latency increases, tasks are offloaded to the ES with more powerful processing capabilities, and there tends to be a preference for a smaller semantic extraction factor, $\mu_{n}$, to maximize the compression rate. This results in a decrease in the accuracy of tasks executed by the ES. Fig. 15 presents the performance of energy consumption and task accuracy under different weights. As the weight of task performance increases, the model allocates more tasks to local, since semantic extraction offloading degrades task performance. However, due to limited local computing capability, longer execution time is required, leading to higher energy consumption. Fig. 16 describes the performance of energy consumption and latency under varying weights. As the weight of latency increases, the model assigns more tasks to the server, leveraging its stronger computational capability, and selects a smaller semantic extraction factor to reduce transmission volume. However, extracting more compact semantic representations results in increased energy consumption.
@@ V-C QoE Preference
V-D Sensitivity Analysis
Figure 17: Sensitivity analysis of logistic slope parameters.
We fix the trained MAPPO policy and vary one of the parameters $\lambda$, $\beta$, and $\eta$ in (10) during evaluation, while keeping the other two unchanged. This allows us to isolate the impact of each parameter on the QoE mapping without introducing policy adaptation effects.
Fig. 17 presents the sensitivity analysis results with respect to $\lambda$, $\beta$, and $\eta$. When the slope of the logistic function is excessively large (or similarly, excessively small), even a slight difference between the latency, energy consumption, or accuracy and their corresponding values under local execution will cause the logistic function to rapidly enter the saturation region. Consequently, the QoE also becomes saturated and loses sensitivity to performance variations.
Based on this observation, we set $\lambda=10^{3}$, $\beta=10$, and $\eta=2$, such that the slope of the logistic function lies within an appropriate range and can effectively capture variations in latency, energy consumption, and accuracy.
VI Conclusion
In this paper, we proposed a new multi-task offloading system and designed a unified quality of experience (QoE) metric to measure overall task offloading performance. Text classification, image classification, and visual question answering (VQA) tasks can be executed locally or offloaded to the edge server after semantic extraction. A semantic-aware multi-agent PPO reinforcement learning algorithm is used to solve the problem of maximizing overall user QoE through the joint optimization of the semantic extraction factor, computing resources, and communication resources. Experimental results show that the proposed method outperforms other benchmarks in overall user QoE. Specifically, the proposed algorithm demonstrates a 12.68% improvement in QoE compared to the semantic-unaware approach. Additionally, the proposed approach can be flexibly adapted to user preferences, fully satisfying latency, energy, and task performance preference requirements. In future work, we intend to develop a robust and general semantic codec that maintains high inference accuracy across heterogeneous data modalities and stochastic, time-varying wireless channels. Besides, we plan to establish rigorous quantitative mathematical relationships between semantic extraction parameters and latency, energy consumption, and task performance, leveraging extensive empirical data.